OpenStudy (anonymous):

x^2 +4x -45=0 by factoring

OpenStudy (anonymous):

What can you multiply to get -45 and add to get 4?

hero (hero):

\(m + n = 4\) \(m \times n = -45\)

hero (hero):

\(9 - 5 = 4\) \((9)(-5) = -45\)

hero (hero):

@stefranc, Do you know what to do next ?